OpenStudy (anonymous):

Which of the following is the simplified form of fifth root of x times the fifth root of x times the fifth root of x times the fifth root of x?

OpenStudy (anonymous):

x to the 1 over fifth power x to the 4 over fifth power x to the 4 over fifth power x

OpenStudy (anonymous):

@hartnn

OpenStudy (cp9454):

i m not getting the question.

OpenStudy (aaronq):

The fifth root of x can be written as \(\sf x^{\frac{1}{5}}\) so your problem says \(\large \sf x^{\frac{1}{5}}*x^{\frac{1}{5}}*x^{\frac{1}{5}}*x^{\frac{1}{5}}=x^{\frac{4}{5}}\)